State and prove weierstrass M- test


1
Expert's answer
2021-03-15T10:29:35-0400

Weierstrass M-test.

 Suppose that (fn) is a sequence of real- or complex-valued functions defined on a set A, and that

there is a sequence of non-negative numbers (Mn) satisfying

"{\displaystyle \forall n\geq 1,\forall x\in A:\ |f_{n}(x)|\leq M_{n},}\\
{\displaystyle \{\displaystyle \sum _{n=1}^{\infty }M_{n}<\infty .}\\
\text{Then the series}\\

{\displaystyle \sum _{n=1}^{\infty }f_{n}(x)}\\
\text{ converges absolutely and uniformly on A}"


Proof


"\\text{Consider the sequence of functions}\\\\\n\n{\\displaystyle S_{n}(x)=\\sum _{k=1}^{n}f_{k}(x).}\\\\\n\n\\text{Since the series} {\\displaystyle \\sum _{n=1}^{\\infty }M_{n}} converges and Mn \u2265 0 \\text{for every} n, \\text{then by the Cauchy criterion},\\\\\n\n{\\displaystyle \\forall \\varepsilon >0:\\exists N:\\forall m>n>N:\\sum _{k=n+1}^{m}M_{k}<\\varepsilon .}\\\\\n\n\\text{For the chosen N},\\\\\n\n{\\displaystyle \\forall x\\in A:\\forall m>n>N}\\\\\n\n{\\displaystyle \\left|S_{m}(x)-S_{n}(x)\\right|=\\left|\\sum _{k=n+1}^{m}f_{k}(x)\\right|{\\overset {(1)}{\\leq }}\\sum _{k=n+1}^{m}|f_{k}(x)|\\leq \\sum _{k=n+1}^{m}M_{k}<\\varepsilon .}\\\\\n\n\\text{(Inequality (1) follows from the triangle inequality.)}\\\\\n\n\\text{The sequence Sn(x) is thus a Cauchy sequence in R or C, and by completeness, it converges to some number S(x) that depends on x. For n\u2009>\u2009N we can write}\\\\\n\n{\\displaystyle \\left|S(x)-S_{n}(x)\\right|=\\left|\\lim _{m\\to \\infty }S_{m}(x)-S_{n}(x)\\right|=\\lim _{m\\to \\infty }\\left|S_{m}(x)-S_{n}(x)\\right|\\leq \\varepsilon .}\\\\\n\n\\text{Since N does not depend on x, this means that the sequence Sn of partial sums converges uniformly to the function S}. \\text{Hence, by definition, the series} {\\displaystyle \\sum _{k=1}^{\\infty }f_{k}(x)} \\text{converges uniformly}.\\\\\n\n\\text{Analogously, one can prove that} {\\displaystyle \\sum _{k=1}^{\\infty }|f_{k}(x)|} \\text{converges uniformly.}"